# Rotating axionic AdS4 black hole dressed with a scalar field

### v1

@@ II Four-dimensional spinning axionic black holes dressed with a scalar field
Figure 1: Left Panel: Graphical representation of ${\ell^{2}f(r)}/{r^{2}}$ as a function of the radial coordinate $r$, with $\eta=B=\ell=1$ for our calculations. Here, Case I (red dotted curve), characterized by $\alpha=1$, indicates the presence of a naked singularity ($04-{\ell^{2}\eta^{2}/\alpha}>$ 0). Case II (black curve), corresponds to the extremal case, given by $\alpha=1/4$. Here, ${\ell^{2}f(r^{*})}/{(r^{*})^{2}}=({\ell^{2}f(r^{*})}/{(r^{*})^{2}})^{\prime}=0$ and this condition emerges when $04-{\ell^{2}\eta^{2}/\alpha}=$ 0. Finally, Case III (dashed blue curve and when $\alpha=1/9$) reveals the situation where $04-{\ell^{2}\eta^{2}/\alpha}<$ 0. Right Panel: Graphical representation of the increasing function ${\ell^{2}f(r)}/{r^{2}}$ at the limit $\alpha\rightarrow 0^{+}$, with $\eta=B=\ell=1$, through the black dotted dashed curve. Here, the location of the event horizon is ensured.
On the other hand, to study the existence of event horizon, we consider the function
@@
${\displaystyle{\left(\frac{\ell^{2}f(r)}{r^{2}}\right)\Big{|}_{r=r^{*}}=1-\frac{\ell^{2}\eta^{2}}{4\alpha}\leq 0,}}$
@@
ensures the existence of the event horizon, as shown in Figure 1 right panel.
Building upon the analysis and framework established for this new four-dimensional axionic spinning BH configuration and the presence of the integration constant $B$ as well as the coupling constants $\alpha$ and $\eta$, in the next section will delve into the thermodynamic quantities.
@@ III Thermodynamics of the rotating black hole solution
The thermodynamic analysis of the previously presented rotating solution (6)-(19) will be conducted using the Euclidean procedure. In this approach, the partition function is equated with the Euclidean path integral, employing the saddle point approximation around the Euclidean continuation of the solution Gibbons:1976ue ; Regge:1974zd . In this Euclidean framework, the Euclidean time $\tau$ is introduced as an imaginary counterpart to the time coordinate $t$, linked through the relation $\tau=it$ from the metric (6). The radial coordinate $r$ spans from the location of, for our case, the outer horizon denoted as $r_{h}$, to +$\infty$+ ∞, while the Euclidean time is made periodic to avoid conical singularities, adhering to the range $0\leq\tau<\beta$, where $\beta=T^{-1}$ and $T$ is the Hawking temperature. With this ansatz, the reduced Euclidean action reads
@@
[…]
@@
The variation with respects the dynamical fields $N$, $f$, $\phi$, $p$, $N^{\theta}$ and $\psi$ yield the following equations
$E_{N}$ ::=:= $\left(1-\frac{\phi^{2}}{6}\right)(rf^{\prime}+f)-\frac{1}{6}r\phi^{\prime}(-rf\phi^{\prime}+r\phi f^{\prime}+4\phi f)-\frac{1}{3}\phi r^{2}f\phi^{\prime\prime}+\Lambda r^{2}+U(\phi)r^{2}$
@@
$E_{f}$ ::=:= ${\frac{1}{6}r\left(2r(\phi^{\prime})^{2}-r\phi\phi^{\prime\prime}\right)N+\frac{1}{6}r\left(r\phi\phi^{\prime}+\phi^{2}-6\right)N^{\prime}=0},$
$E_{\phi}$ ::=:= ${-\frac{1}{3}\phi r^{2}fN^{\prime\prime}-\frac{1}{6}r\left(6rf\phi^{\prime}+4\phi f+3\phi rf^{\prime}\right)N^{\prime}-\frac{1}{6}\Big{(}4\phi rf^{\prime}+2\phi f+r^{2}\phi f^{\prime\prime}+6\phi^{\prime}r^{2}f^{\prime}}$
@@
$E_{p}$ ::=:= $12Np+r^{4}(\phi^{2}-6)(N^{\theta})^{\prime}=0,$
$E_{N^{\theta}}$ ::=:= $p^{\prime}=0,$
$E_{\varphi}$ ::=:= $\partial_{\varphi}^{2}\psi=0.$
@@
$\delta B_{E}$ $=$ $\beta\sigma\Big{[}N\left(\frac{1}{6}r^{2}\phi\phi^{\prime}+r\left(\frac{1}{6}\phi^{2}-1\right)\right)\delta f-\frac{1}{6}r^{2}\left(N\left(4f\phi^{\prime}+\phi f^{\prime}\right)+2N^{\prime}\phi f\right)\delta\phi$
$+$ $\frac{N}{3}\phi r^{2}f\delta(\phi^{\prime})-N^{\theta}dp\Big{]}_{r_{h}}^{+\infty}-\beta\int_{r_{h}}^{+\infty}N\varepsilon dr\left[\int_{0}^{\sigma_{\theta}}d\theta\partial_{\varphi}\psi\delta\psi\right]_{0}^{\sigma_{\varphi}},$
@@
[…]
Now we are in a position to compute the boundary term $B_{E}$. At the infinity, from eqs. (12)-(13) we have that
$\delta N^{\theta}\big{|}_{+\infty}$ $=$ ${\frac{\sqrt{\alpha}{B}^{2}dB}{6{r}^{3}}}+{\frac{\sqrt{\alpha}{B}^{4}dB}{6{r}^{5}}}+O\left(\frac{1}{r^{6}}\right),$
$\delta f\big{|}_{+\infty}$ $=$ ${-\frac{\eta B^{2}dB}{6r}-\frac{\eta B^{4}dB}{6r^{3}}+O\left(\frac{1}{r^{4}}\right),}$ (32)
while that with respect to $p$, and the scalar fields ($\phi,\psi$):
$\delta\phi\big{|}_{+\infty}$ $=$ $\frac{\sqrt{6}dB}{r}-\frac{2\sqrt{6}BdB}{r^{2}}+O\left(\frac{1}{r^{3}}\right),$
$\delta\psi\big{|}_{+\infty}$ $=$ $\varphi\delta B,\qquad\delta p\big{|}_{+\infty}=\delta\left(-\frac{1}{12}{\sqrt{\alpha}B^{3}}\right).$ (33)
@@
${\delta B_{E}\big{|}_{+\infty}=\frac{\beta\sigma\eta B^{2}dB}{6}\Rightarrow B_{E}\big{|}_{+\infty}=\frac{\beta\sigma\eta B^{3}}{18}.}$
@@
we have
$\delta B_{E}\big{|}_{r_{h}}=\delta\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)+\beta\Omega\delta(\sigma p)+\beta\Psi\delta(\sigma B)\Rightarrow$
$B_{E}\big{|}_{r_{h}}=\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)+\beta\Omega(\sigma p)+\beta\Psi(\sigma B),$
@@
[…]
@@
$B_{E}$ $=$ ${B_{E}\big{|}_{+\infty}-B_{E}\big{|}_{r_{h}}=\frac{\beta\sigma\eta B^{3}}{18}-\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)-\beta\Omega(\sigma p)-\beta\Psi_{a}(\sigma B).}$ (37)
The identification of the thermodynamic parameters is achieved recalling that the Euclidean action is related to the Gibbs free energy $G$ as
@@
where ${\cal{M}}$ is the mass, ${\cal{S}}$ the entropy and, as before, $\Omega$ (respectively $\Psi_{a}$) is the chemical potential associated with the angular momentum ${\cal{J}}$ (respectively axionic charge $\mathcal{\omega}_{a}$). Hence, comparing (37) with (38) one obtains that the entropy reads
${\cal{S}}=\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}},$ (39)
while the mass ${\cal{M}}$, angular momentum ${\cal{J}}$, angular velocity $\Omega$, axionic potential $\Psi_{a}$, and axionic charge ${\cal{\omega}}_{a}$ are given by
@@
It is reassuring to check that with the Hawking temperature,
$T=\frac{N(r)f^{\prime}(r)}{4\pi}\Big{|}_{r=r_{h}}={\frac{(N^{\theta}(r))^{\prime}(2\sqrt{\alpha}N^{\theta}(r)-\eta)r^{2}}{4\pi\sqrt{\alpha}}\Big{|}_{r=r_{h}},}$ (44)
the first law
${\delta{\cal{M}}=T\delta{\cal{S}}+\Omega\delta{\cal{J}}+\Psi_{a}\delta{\omega_{a}}},$ (45)
is satisfied.
For the non-spinning situation, represented via the line element (20) with $N=1$, together with the metric function $f$ (21), the scalar field $\phi$ (14) and the potential $U(\phi)$ (16) with $U_{\alpha}(\phi)=0$, we note that the entropy ${\cal{S}}$ as well as the mass ${\cal{M}}$ take the form given in (39) and (40) respectively, while that the Hawking Temperature $T$ reads
@@
and the first law $\delta{\cal{M}}=T\delta{\cal{S}}$ holds.

### v2

@@ II Four-dimensional spinning axionic black holes dressed with a scalar field
Figura 1: Left Panel: Graphical representation of ${\ell^{2}f(r)}/{r^{2}}$ as a function of the radial coordinate $r$, with $\eta=B=\ell=1$ for our calculations. Here, Case I (red dotted curve), characterized by $\alpha=1$, indicates the presence of a naked singularity ($4-{\ell^{2}\eta^{2}/\alpha}>0$). Case II (black curve) corresponds to the extremal case, given by $\alpha=1/4$. Here, ${\ell^{2}f(r^{*})}/{(r^{*})^{2}}=({\ell^{2}f(r^{*})}/{(r^{*})^{2}})^{\prime}=0$ and this condition emerges when $4-{\ell^{2}\eta^{2}/\alpha}=0$. Finally, Case III (dashed blue curve and when $\alpha=1/9$) reveals the situation where $4-{\ell^{2}\eta^{2}/\alpha}<0$. Right Panel: Graphical representation of the increasing function ${\ell^{2}f(r)}/{r^{2}}$ at the limit $\alpha\rightarrow 0^{+}$, with $\eta=B=\ell=1$, through the black dotted dashed curve. Here, the location of the event horizon is ensured.
On the other hand, to study the existence of the event horizon (or outer), we consider the function
@@
${\left(\frac{\ell^{2}f(r)}{r^{2}}\right)\Big|_{r=r^{*}}=1-\frac{\ell^{2}\eta^{2}}{4\alpha}\leq 0,}$}
@@
ensures the existence of an event horizon, as shown in Figure 1 right panel.
Building upon the analysis and framework established for this new four-dimensional axionic spinning BH configuration and the presence of the integration constant $B$ as well as the parameter constants $\alpha$ and $\eta$, in the next section, we will delve into the thermodynamic quantities.
@@ III Thermodynamics of the rotating black hole solution
The thermodynamic analysis of the previously presented rotating solution (6)-(23) will be conducted using the Euclidean procedure. In this approach, the partition function is equated with the Euclidean path integral, employing the saddle point approximation around the Euclidean continuation of the solution Gibbons:1976ue ; Regge:1974zd . In this Euclidean framework, the Euclidean time $\tau$ is introduced as an imaginary counterpart to the time coordinate $t$, linked through the relation $\tau=it$ from the metric (6). The radial coordinate $r$ spans from the location of, for our case, the outer horizon denoted as $r_{h}$, to $\infty$, while the Euclidean time is made periodic to avoid conical singularities, adhering to the range $0\leq\tau<\beta$, where $\beta=T^{-1}$ and $T$ is the Hawking temperature. Following this standard minisuperspace procedure, once the geometric symmetries are imposed, the ansatz defines a consistent truncation of the full configuration space. Therefore, the functions are treated as independent fields of a one-dimensional effective action, and the reduced Euclidean action reads:
@@
$+$ $\frac{1}{2}\beta\int_{r_{h}}^{+\infty}dr\int_{0}^{\sigma_{\theta}}d\theta\int_{0}^{\sigma_{\varphi}}d\varphi\big({N}\varepsilon(\phi)\left(\partial_{\varphi}\psi\right)^{2}\big)+B_{E},$
@@
The variation with respect to the dynamical fields $N$, $f$, $\phi$, $p$, $N^{\theta}$ and $\psi$ yield the following equations
$E_{N}$ $:=$ $\left(1-\frac{\phi^{2}}{6}\right)(rf^{\prime}+f)-\frac{1}{6}r\phi^{\prime}(-rf\phi^{\prime}+r\phi f^{\prime}+4\phi f)-\frac{1}{3}\phi r^{2}f\phi^{\prime\prime}+\Lambda r^{2}+U(\phi)r^{2}$
@@
$E_{f}$ $:=$ ${\frac{1}{6}r\left(2r(\phi^{\prime})^{2}-r\phi\phi^{\prime\prime}\right)N+\frac{1}{6}r\left(r\phi\phi^{\prime}+\phi^{2}-6\right)N^{\prime}=0},$
$E_{\phi}$ $:=$ ${-\frac{1}{3}\phi r^{2}fN^{\prime\prime}-\frac{1}{6}r\left(6rf\phi^{\prime}+4\phi f+3\phi rf^{\prime}\right)N^{\prime}-\frac{1}{6}\Big(4\phi rf^{\prime}+2\phi f+r^{2}\phi f^{\prime\prime}+6\phi^{\prime}r^{2}f^{\prime}}$
@@
$E_{p}$ $:=$ $12Np+r^{4}(\phi^{2}-6)(N^{\theta})^{\prime}=0,$
$E_{N^{\theta}}$ $:=$ $p^{\prime}=0,$
$E_{\varphi}$ $:=$ $\partial_{\varphi}^{2}\psi=0.$
@@
$\delta B_{E}$ $=$ $\beta\sigma\Big[N\left(\frac{1}{6}r^{2}\phi\phi^{\prime}+r\left(\frac{1}{6}\phi^{2}-1\right)\right)\delta f-\frac{1}{6}r^{2}\left(N\left(4f\phi^{\prime}+\phi f^{\prime}\right)+2N^{\prime}\phi f\right)\delta\phi$ (36)
$+$ $\frac{N}{3}\phi r^{2}f\delta(\phi^{\prime})-N^{\theta}dp\Big]_{r_{h}}^{+\infty}-{\beta\int_{r_{h}}^{+\infty}drN\varepsilon\left[\int_{0}^{\sigma_{\theta}}d\theta\partial_{\varphi}\psi\delta\psi\right]_{0}^{\sigma_{\varphi}}},$
@@
$\sigma=\int_{0}^{\sigma_{\theta}}d\theta\int_{0}^{\sigma_{\varphi}}d\varphi=\sigma_{\theta}\sigma_{\varphi}.$
Here, we emphasize that the Euclidean action used in our thermodynamic analysis is not based solely on the Ricci scalar $R$, but rather includes the full contributions from the non-minimally coupled scalar field and the axionic sector. Additionally, the Euclidean action (34) is constructed by evaluating the complete Lagrangian and including a boundary term $B_{E}$ whose variation $\delta B_{E}$ is derived explicitly to ensure that $\delta I_{E}=0$ on-shell (see eq. (36)). This boundary analysis is crucial and ensures that the thermodynamic quantities obtained via the saddle-point approximation are physically meaningful and consistent with the first law.
Now we are in a position to compute the boundary term $B_{E}$. For this, we calculate the variation of the dynamical fields $\delta N^{\theta}$, $\delta f$, $\delta\psi$, $\delta\phi$, and $\delta p$, where the constant parameters $\alpha$ and $\eta$ are fixed. The variations are performed with respect to the integration constant $B$ as well as the location of the event horizon $r_{h}$. At infinity, we have that
$\delta N^{\theta}\big|_{+\infty}$ $=$ ${{\frac{\sqrt{\alpha}{B}^{2}\delta B}{6{r}^{3}}}+{\frac{\sqrt{\alpha}{B}^{4}\delta B}{6{r}^{5}}}+O\left(\frac{1}{r^{6}}\right)},$
$\delta f\big|_{+\infty}$ $=$ ${{-\frac{\eta B^{2}\delta B}{6r}-\frac{\eta B^{4}\delta B}{6r^{3}}+O\left(\frac{1}{r^{4}}\right),}}$ (37)
whereas with respect to $p$, and the scalar fields ($\phi,\psi$):
$\delta\phi\big|_{+\infty}$ $=$ ${\frac{\sqrt{6}\delta B}{r}-\frac{2\sqrt{6}B\delta B}{r^{2}}+O\left(\frac{1}{r^{3}}\right)},$
$\delta\psi\big|_{+\infty}$ $=$ $\varphi\delta B,\qquad\delta p\big|_{+\infty}={-\frac{\sqrt{\alpha}B^{2}\delta B}{4}=\delta\left(-\frac{1}{12}{\sqrt{\alpha}B^{3}}\right).}$ (38)
@@
$\delta B_{E}\big|_{+\infty}={\frac{\beta\sigma\eta B^{2}\delta B}{6}=\delta\left(\frac{\beta\sigma\eta B^{3}}{18}\right)}\Rightarrow B_{E}\big|_{+\infty}=\frac{\beta\sigma\eta B^{3}}{18}.$}
@@
we have that the last expression of eq. (36) (with $N=1$) takes the form:
$\int_{r_{h}}^{+\infty}dr\varepsilon\left[\int_{0}^{\sigma_{\theta}}d\theta\partial_{\varphi}\psi\delta\psi\right]_{0}^{\sigma_{\varphi}}=\int_{r_{h}}^{+\infty}dr(B\varepsilon)(\sigma\delta B)=\Psi_{a}\delta(\sigma B)$
and
$\delta B_{E}\big|_{r_{h}}=\delta\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(r_{h}+B)^{2}}\right)+\beta\Omega\delta(\sigma p)+\beta\Psi_{a}\delta(\sigma B)\Rightarrow$
$B_{E}\big|_{r_{h}}=\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(r_{h}+B)^{2}}\right)+\beta\Omega(\sigma p)+\beta\Psi_{a}(\sigma B),$
@@
${\Psi_{a}}$ $=$ $\int_{r_{h}}^{+\infty}{drB\varepsilon\big(\phi(r)\big)=\left[-\frac{B^{3}\alpha(r+B)(-r^{2}+2Br+2B^{2})}{1152r^{4}}-\frac{B^{2}\alpha}{2304}\ln\left(\frac{2B+r}{r}\right)\right]_{r_{h}}^{+\infty}}$ (41)
@@
$B_{E}$ $=$ ${B_{E}\big|_{+\infty}-B_{E}\big|_{r_{h}}=\frac{\beta\sigma\eta B^{3}}{18}-\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(r_{h}+B)^{2}}\right)-\beta\Omega(\sigma p)-\beta\Psi_{a}(\sigma B).}$ (42)
The identification of the thermodynamic parameters is achieved by recalling that the Euclidean action is related to the Gibbs free energy $G$ as
@@
[…]
${\cal{S}}=2\pi\sigma r_{h}^{2}\left(1-\frac{1}{6}\phi(r_{h})^{2}\right)=\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(r_{h}+B)^{2}},$ (44)
[…]
@@
The Hawking temperature reads
$T=\frac{\kappa}{2\pi}=\frac{N(r)f^{\prime}(r)}{4\pi}\Big|_{r=r_{h}}={\frac{(N^{\theta}(r))^{\prime}(2\sqrt{\alpha}N^{\theta}(r)-\eta)r^{2}}{4\pi\sqrt{\alpha}}\Big|_{r=r_{h}},}$ (49)
where $\kappa$ is the surface gravity, given by
$\kappa=\sqrt{-\frac{1}{2}\left(\nabla_{\mu}\xi_{\nu}\right)\left(\nabla^{\mu}\xi^{\nu}\right)},$ (50)
with a Killing vector $\xi^{\mu}=(\partial_{t})^{\mu}-\Omega(\partial_{\theta})^{\mu}$ and $\Omega=-N^{\theta}(r_{h})$. Here, it is important to note that:
$\delta{\cal{M}}$ $=$ $\frac{\sigma\eta B^{2}\delta B}{6},$
$T\delta{\cal{S}}$ $=$ $T(\partial_{r_{h}}{\cal{S}})\delta r_{h}+T(\partial_{B}{\cal{S}})\delta B=\frac{\sigma\eta B^{2}\delta B}{6}-{\frac{\sigma N^{\theta}(r_{h})}{3}{\sqrt{\alpha}B^{2}}}\delta B,$
$\Omega\delta{\cal{J}}$ $=$ ${\frac{\sigma N^{\theta}(r_{h})}{4}{\sqrt{\alpha}B^{2}}}\delta B,$
$\Psi_{a}\delta{\omega_{a}}$ $=$ ${\frac{\sigma N^{\theta}(r_{h})}{12}{\sqrt{\alpha}B^{2}}}\delta B,$
ensuring the first law
${\delta{\cal{M}}=T\delta{\cal{S}}+\Omega\delta{\cal{J}}+\Psi_{a}\delta{\omega_{a}}}.$ (51)
Here, we are using the fact that from eqs. (16) and (19):
$1-\frac{{\ell^{2}}\eta}{\sqrt{\alpha}}N^{\theta}(r_{h},B)+\ell^{2}(N^{\theta}(r_{h},B))^{2}=0,$
and from this last expression:
$\left(\partial_{r_{h}}N^{\theta}\right)\delta r_{h}+\left(\partial_{B}N^{\theta}\right)\delta B=0,$
with $\partial_{a}{\cal{S}}:=\partial{\cal{S}}/\partial a$ and $\partial_{a}N^{\theta}:=\partial N^{\theta}/\partial a$. To reinforce the above, the entropy $\cal{S}$} (eq. (44)) can be recovered using the Wald formula Wald:1993nt ; Iyer:1994ys , denoted as ${\cal{S}}_{W}$, where
${\cal{S}}_{W}=-2\pi\oint_{\Sigma}d^{2}x\sqrt{|\gamma|}P^{\mu\nu\sigma\rho}\,\varepsilon_{\mu\nu}\,\varepsilon_{\sigma\rho}=2\pi\sigma r_{h}^{2}\left(1-\frac{1}{6}\phi(r_{h})^{2}\right)=\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(r_{h}+B)^{2}},$
and
$P^{\mu\nu\sigma\rho}=\frac{\partial(\mathcal{L}_{g}+\mathcal{L}_{\phi})}{\partial R_{\mu\nu\sigma\rho}}=\frac{1}{4}\left(1-\frac{1}{6}\phi^{2}\right)\left(g^{\mu\sigma}g^{\nu\rho}-g^{\mu\rho}g^{\nu\sigma}\right).$
Here, the expression is evaluated at the spatial section $\Sigma$ of the event horizon. The quantity $|\gamma|$ represents the determinant of the induced metric on $\Sigma$, while $\varepsilon_{\mu\nu}$ denotes the binormal vector
$\varepsilon_{\mu\nu}=-\varepsilon_{\nu\mu}:=\frac{1}{\kappa}\,\nabla_{\mu}\xi_{\nu},$
where $\kappa$ is the surface gravity shown previously in eq. (50). The mass $\mathcal{M}$ and angular momentum $\mathcal{J}$ can be recovered via the Brown-York method Brown:1992br ; Brown:1994gs ; Balasubramanian:1999re , where the explicit steps are given in Appendix B.
[…]
@@
and the first law $\delta{\cal{M}}=T\delta{\cal{S}}$ holds, where now $r_{h}$ corresponds to the location of the event horizon for the non-spinning configuration.